# Dynamical and structural properties of an absorbing phase transition: a case study from granular systems

### v1

@@ II.2 The coarse-grained model
Event driven molecular dynamics simulations [94] of this 2D model are performed with up to $1.5\times 10^{7}$ particles in a square box with periodic boundary conditions. We define an arbitrary unit of time $\hat{\tau}$ and let the system evolve over a time $t$ large enough for a steady state temperature $T^{ss}$ to be reached. The system-wide granular temperature $T$ of the 2D model, defined as in Eq. (1), will play the role of a mean field order parameter and is the equivalent of the $xy$ temperature defined for the realistic model. We define again the steady state temperature $T^{ss}$ as the temperature measured after the system has relaxed. As shown in Fig. 3c, we observe that by setting $01/\tau_{s}=$ 0 (i.e. no synchronization, circular plot markers) the system undergoes a continuous transition as the packing fraction of particles is varied. However, once synchronization is introduced by setting $01/\tau_{s}>$ 0 a discontinuous transition is observed, in agreement with the results found in the realistic model. For fixed values of $\Delta$ and $\alpha$, two timescales are at play, as seen in Fig. 3d. On the one hand, for a fixed value of $\tau_{s}$, the critical packing fraction increases with increasing drag coefficient $\gamma$. Indeed, higher drag implies that a denser system is needed to ensure sufficiently frequent collisions to sustain an active steady state. On the other hand, at a fixed value of $\gamma$, the critical packing fraction increases upon decreasing $\tau_{s}$, as this makes it easier to synchronize and hence lose energy. Another way to look at this is through the frequency of collision. When synchronization is at play in our system, we expect that at the critical packing fraction, the rate at which a particle synchronizes $1/\tau_{s}$ is of the same order of the frequency of collision which is itself, at a given temperature, an increasing function of $\phi$.
Figure 4: a, b, c) Power law behavior of observables for the realistic model with parameters $A=0.085\sigma$, $h=1.5077\sigma$ and $N=30000$ leading to $\phi_{c}\simeq 0.186$. d, e, f) Power law behavior of observables for the coarse-grained model with parameters $\Delta/(\sigma\hat{\tau})=0.0015$, $\gamma\hat{\tau}=0.01$, $\alpha=0.95$, $N=10^{6}$ leading to $\phi_{c}\simeq 0.1509$. The dashed lines represent the power law expected from a $2d$ system belonging to the conserved directed percolation universality class [19, 95]. a, d) Order parameter as a function of $\phi-\phi_{c}$. Expected exponent $\beta\simeq 0.64$ with $T^{ss}\sim(\phi-\phi_{c})^{\beta}$. b, e) Average over 100 runs of the evolution of the order parameter at $\phi_{c}$. Expected exponent $a\simeq 0.42$ with $T(t)\sim t^{-a}$. c, f) Time to relax to the steady state as a function of $|\phi-\phi_{c}|$. Expected exponent $\gamma\simeq 1.3$ with $\tau_{r}\sim|\phi-\phi_{c}|^{-\gamma}$.
@@ III.1 Nature of the transitions
Thus far, we have discussed the APT in terms of a discontinuous or continuous transition, simply based on whether there is an obvious jump in kinetic energy as a function of packing fraction. In this section, we take a closer look at the behavior around the transition point in order to examine the nature of the transition in more detail, for both the realistic and coarse-grained model.
@@
For continuous transitions, in systems characterized by an infinite number of absorbing states and a conserved number of particles without additional symmetries, the associated phase transition is believed to belong to the conserved directed percolation universality class [96, 29, 97, 93]. Evidence of this universality class, from the analysis of power law scaling, is given in Fig. 4 where simulation results for both the realistic and the coarse-grained models are compared to the expected result for systems belonging to the conserved directed percolation universality class. In particular, panels a), b) and c) present data of the realistic model while d), e) and f) present those of the coarse-grained model. In panels a) and d), we show that the order parameter variation, as a function of the distance from the critical point, obeys a power-law scaling characterized by an exponent very close to what is expected for conserved directed percolation universality class. In panels (b) and (e), we give the time evolution of the order parameter at the critical packing fraction $\phi_{c}$. For the coarse-grained model, the expected scaling closely matches the evolution of the order parameter. For the realistic model, however, it is more challenging to discern the correct scaling due to the limited simulation time and the difficulty to accurately pinpoint the transition point. In panels d) and f), we provide data for the characteristic time $\tau_{r}$ to reach the steady state. Above $\phi_{c}$, $\tau_{r}$ is defined as the time at which the system first reaches its steady-state value within an arbitrary precision, which we set to 10 percent of the average steady-state temperature. These curves match well the expected scaling. Below $\phi_{c}$, $\tau_{r}$ is defined as the first time the system reaches an arbitrarily small value. We find significant variations in the curves depending on the chosen threshold, since the energy decreases as an exponential after the last collision. Here, we define the threshold as the lowest metastable steady-state temperature observed in a system that eventually decayed to the absorbing state. Note that the measured exponent is known to be dependent on the initial conditions [95]. However, from the overall agreement with the expected scaling laws, we believe that we have good evidences to conclude that the phase transition belongs to the conserved directed percolation universality class.
@@
It is worth noting that our focus has been on systems exhibiting their phase transition at relatively high critical packing fraction ($\phi_{c}>0.1$). In a very similar system [93], it has been found that if the transition occurs at a very small packing fraction, the critical exponents of the system align with those of the mean-field conserved directed percolation universality class. This can be understood from the fact that, at criticality, the typical interaction length of the system is given by the mean free path. At low $\phi_{c}$, Lei and Ni [93] argued that the system is effectively governed by long range interactions, explaining why mean field exponents are found. A study of the crossover between the short and long range effective interaction for similar transitions can be found in Ref. 2.
@@
Figure 5: Comparison of the order parameter evolution as a function of packing fraction in the realistic a) and coarse-grained model b) for different system sizes. The main figures represent the evolution of the steady state temperature in the discontinuous case (we recall that the temperature is the $xy$-kinetic energy for the realistic model and the usual kinetic energy for the coarse-grained model) for different system sizes and insets show the evolution of the order parameter of the last (the most dilute) active steady state ($\Delta T_{ss}$) as a function of the system size a) realistic model: $A=0.0626\sigma$ and $h=1.95\sigma$. b) Coarse-grained model: $\hat{\tau}\Delta/\sigma=0.015$, $\alpha=0.95$, $\hat{\tau}\gamma=0.01$ and $\tau_{s}/\hat{\tau}=6$.
@@
Having clarified the nature of the continuous transition, we now turn to the discontinuous one and perform a finite size analysis to prove that the finite discontinuity does persist in the infinite system size limit. We define $\Delta T_{ss}$ which corresponds to the observed value of the order parameter of the most dilute system in the active state at a given system size and simulation time. For a continuous transition, we expect that at finite times, $\Delta T_{ss}$ decreases with $N$ and reaches 0 at infinite system size since the order parameter must vanish continuously. On the contrary, for discontinuous transition, $\Delta T_{ss}$ should not vanish. Evolution of the order parameter as a function of $\phi$ close to the discontinuous transition as well as $\Delta T_{ss}$ as a function of system size in inset are given, for the realistic and coarse-grained models in the insets of Fig 5 a) and b) respectively. Starting from small system sizes, we see that increasing its size allows us to obtain a transition at a lower density, as would be expected as well for a continuous transition. However, beyond a given system size, $\Delta T_{ss}$ ceases to decrease and instead increases with size, indicating a discontinuous jump in the order parameter in the thermodynamic limit and confirming the first-order nature of the transition.
@@
We interpret such non-monotonic behavior of $\Delta T_{ss}$ as a nucleation dynamics better understood directly from the average time necessary for the system to reach an absorbing state at a given packing fraction in the metastable zone. We denote this time $t_{\mathrm{wait}}$ and study its behavior as a function of the number of particles for the coarse-grained model in Fig. 6. The time $t_{\mathrm{wait}}$ acts as a proxy for the nucleation time, since the growth of the nucleus after it has reached its critical size should be fast compared to $t_{\mathrm{wait}}$. We only perform the analysis for the coarse-grained model due to the prohibitive simulation time of the realistic model. For small system sizes, the waiting time increases with $N$ since the fluctuation required for the system to reach an absorbing state is of the order of the size of the system. Then, the waiting time reaches a maximum, which corresponds to a system size close to the size of the critical nucleus. It then decays due to the increase of available nucleation sites, until reaching a constant at larger system sizes due to the finite cooling rate imposed by $\gamma$ and the frequency of collision. Note that the decrease of the waiting time for system sizes larger than the critical nucleus is very sharp, while in an equilibrium we would expect a decay proportional to $1/N$. This might be the result of hydrodynamical instabilities or non-equilibrium structural properties at small $k$, inaccessible to small system sizes, which are common in granular systems [98, 99, 100]. Note also that contrary to equilibrium systems, the nucleation dynamics, and especially the growth of the nucleus, exhibit an exotic behavior, both for the coarse-grained and the realistic model. Since the order parameter is the activity or the temperature of the system, the nucleus cannot be stationary because, during its growth, active particles around it will quickly convert its inactive grains into active ones. We instead observe damped traveling waves of energy and density, slowly decreasing in amplitude until the absorbing state is reached (see Video 1 for the Realistic model and Video 2 for the coarse grained model). Moreover, contrary to Ref. 31 we do not find a strongly diverging radius of the critical nucleus as we approach the stable fluid region but merely a strong increase of the effective energy barrier of nucleation. In the inset of Fig. 6 we confirm that the statistics of the waiting time are Poissonian when the reaching of the absorbing state is controlled by nucleation [101].
Interestingly, a very similar system has been studied by Lei, Hu, and Ni [102, 103]. While it does not contain synchronization, the collision dynamics is effectively similar to ours. In their model, if two particles collide with an energy smaller than some threshold, then the particles undergo elastic collision, otherwise they gain a finite amount of energy. A viscous drag is also applied during the free flight. In our model, particles that do not have a lot of energy are, on average, those which traveled for a long time and are slowed down due to drag. Hence, they are likely to be synchronized and will therefore undergo collisions with $\Delta=0$. This might mimic the effect of the energy-activated active collision of the models studied by Lei et al. Thus, the rough phenomenology of both models is similar even though Refs. 102, 103 do not consider dissipative collisions, and drag is de facto their only source of dissipation. Interestingly, they find that the pathway of the metastable fluid to the absorbing state is seemingly not driven by a nucleation process but by large wavelength fluctuations which are suppressed due to the non-equilibrium nature of the system. Indeed, they find that the metastable fluid has suppressed density fluctuations. On this basis, they argue that the non-equilibrium equivalent of the equilibrium metastable state becomes kinetically stable in the thermodynamic limit, due to the suppression of large wavelength fluctuations. The shrinking of the metastable region with system size is also observed using the minimal field theory of the conserved directed percolation universality class, modified to obtain a first order transition [51, 104]. This is at odds with our results, as well as what was found in Ref 31. We believe the discrepancy is caused by the presence of dissipative collisions in our system that induces the formation of clusters, which enhance long wavelength fluctuations and allow for the regular nucleation pathway to the transition. We provide additional discussions on this matter in Appendix. C.
@@
Figure 6: Waiting time in the coarse-grained model for a system initialized in an active state to reach an absorbing state for different densities in the metastable zone; $\alpha=0.95$, $\hat{\tau}\gamma=0.01$, $\hat{\tau}/\tau_{s}=1/6$ and $\hat{\tau}\Delta/\sigma=0.015$. The inset shows the average waiting time as a function of its standard deviation. The dashed line is a linear scaling, showing that the process is Poissonian. Each point corresponds to an average over 50 independent runs.
@@ III.2 Structural properties of the transitions
Figure 7: Zoomed snapshots of the systems above ($\phi_{c}^{+}$, left) and below ($\phi_{c}^{-}$, right) the critical packing fraction for the realistic and coarse-grained model. The particles with small velocities are drawn in black. The insets above $\phi_{c}$ represent the velocity distribution of particles corresponding to the snapshots in units of $\Delta$ for the coarse-grained model and $\sigma f$ for the realistic model. The colors of the histograms of velocities match those of the particles in the snapshots (inactive are in black and active are in orange). The other insets represent the structure factor $\bm{S}_{\rho\rho}(k)$ of the corresponding snapshots as a function of wavenumber in units of $1/\sigma$.
Having clarified the nature of the transition, we now focus on dynamical and structural properties of the system close to it. In Fig. 7, we present, for both the coarse-grained model and the realistic one, snapshots above and below $\phi_{c}$ for both type of transition, and the probability distribution of the velocities in the diffusive phase, and the long-wavelength behavior of the structure factor in the absorbing phase.
In the case of a continuous transition, the system reaches the absorbing state due to viscous drag forces. As commonly observed in similar situations, the critical point is characterized by avalanches of active particles that span the whole system [105, 106, 107]. This behavior is exemplified in the snapshots and leads to a distinctly identifiable population of inactive particles, discernible through the peak observed in the velocity distribution at the origin.
@@
Close to the transition point, $S_{\rho\rho}$ exhibits a scaling behavior of the form $S_{\rho\rho}(|\bm{k}|\rightarrow 0)\sim k^{\nu}$, with $\nu\simeq 0.5$ [108]. The vanishing of the structure factor at $k=0$ is called hyperuniformity [109, 110, 111] and is directly related to a suppression of density fluctuations on large length scales. This trend is consistent with the typical hyperuniform behavior observed at the critical point associated with the conserved directed percolation universality class [112, 93]. It has been observed that this scaling is expected to hold exactly at the transition point and to be lost below or above it [113]. In our models, however, system sizes are presumably not large enough for this hyperuniform scaling to break. Very recent works have successfully predicted hyperuniform scaling of the structure factor from a Renormalization Group approach by mapping the conserved directed percolation problem to the interface position at depinning for an elastic manifold [108]. A direct analysis of the hyperuniformity from the Renormalization Group of the conserved directed percolation field theory was also performed in Ref. [114] at first loop.
@@
Finally, we report new intriguing hyperuniform scaling for the longitudinal and transverse velocity fields at criticality in Appendix. B. The velocity distributions in the active state exhibit peculiar scaling. As already noted, a finite portion of the system is almost completely inactive, hence there is a sharp peak at $v=0$, both for the realistic model and the coarse-grained model to the velocity distribution. Moreover, at small velocity, we observe an exponential distribution with a nontrivial power: $\exp(-a|v|^{\delta})$ with $\delta\leq 1$, which then morphs into a simple exponential scaling: $\exp(-b|v|)$ at $v\simeq\Delta$. This effect is not surprising as non-Gaussian scalings are common in granular systems and, notably, it is known that the homogeneous cooling state exhibits a simple exponential tail while the homogeneously heated granular gas has an exponential tail with exponent $\delta=3/2$ [115, 116, 117]. In particular, a simple exponential tail appears frequently [118, 119] and is conjectured to be universal for systems with adiabatic energy change. However, this is not the case in our system, since the energy change during a collision is large compared to the energy of the particles. In our case, the pronounced non-Gaussian behavior is most likely due to the finite jump processes affecting the velocities, effectively transforming the dynamics into a generalized Ornstein-Uhlenbeck process with discrete jumps of size ±$\pm\Delta$. This makes the system analogous to a stochastic resetting problem, where an overdamped particle in a harmonic potential is reset away from the origin. Such processes are known to asymptotically yield stationary distributions with exponential or power-law tails [120, 121].In Appendix A, we provide a potential explanation of this peculiar velocity distribution using an approximation for the Boltzmann equation.
@@
In the discontinuous case, the system primarily transitions into the absorbing phase in a manner similar to the cooling behavior observed in a granular gas. This is because most of the energy is dissipated through collisions between synchronized particles, with the drag playing a secondary role in this process. Interestingly, the presence of viscous friction is not a prerequisite for observing a transition. While it helps the system to reach sooner an absorbing state, the dissipative character of synchronized collisions alone is sufficient to deplete the entire energy of the system in the infinite time limit. This implies the formation of clusters similar to those observed in the inhomogeneous cooling state [122, 98, 123] for the realistic model, as can be seen in the snapshots and from the upturn of the structure factor at $k\rightarrow 0$. A scaling close to $k^{-3}$, but less steep is observed in the intermediate wavenumber region, which is roughly consistent with Porod’s law [124, 125]. The inhomogeneous cooling is however arrested at a finite time due to the drag. In the coarse-grained model, as collisions lead to a complete asynchronization of particles, the formation of clusters due to dissipative collisions is less prominent because high density regions of synchronized particles will quickly be shattered by any collision. However, even if the growth of the nucleus cannot proceed as the beginning of the inhomogeneous cooling in the coarse-grained model, Porod’s law is approximately verified.
@@
The active state of the discontinuous transition does not exhibit a population of inactive particles but still has a marked non-Gaussian behavior, that, contrary to the case of the continuous transition, can be in part attributed to the dissipative collisions. We also provide the structure factor of the fluid close to the metastable region, which is in itself interesting. Lei and Ni [103] observed a hyperuniform scaling $S\sim k^{1.2}$ in a metastable fluid at very small wavenumbers. As already mentioned before, due to the dissipative collisions between synchronized particles, we instead observe a peak in the structure factor corresponding to the inverse average cluster size. This peak might be followed by a hyperuniform scaling at lower $k$. Verifying this hypothesis is however complicated due to the large system sizes required. This peak disappears with the disappearance of synchronized particles (not shown). We show in Appendix C that without dissipative collisions ($\alpha=1$) close to the metastable fluid, we recover a structure factor similar to the one found by Lei and Ni.

### v2

@@ II.2 The effective 2D model
Event driven molecular dynamics simulations [smallenburg2022efficient] of this 2D model are performed with up to $1.5\times 10^{7}$ particles in a square box with periodic boundary conditions. We ensure that simulations are run long enough to let the system reach a steady-state temperature $T^{ss}$. The system-wide granular temperature $T$ of the 2D model, defined as in Eq. (1), will play the role of a mean field order parameter and is the equivalent of the $xy$ temperature defined for the realistic quasi-2D model. We define again the steady state temperature $T^{ss}$ as the temperature measured after the system has relaxed. As shown in Fig. 3c, we observe that by setting $1/\tau_{s}=0$ (i.e. no synchronization, circular plot markers) the system undergoes a continuous transition as the packing fraction of particles is varied. However, once synchronization is introduced by setting $1/\tau_{s}>0$ a discontinuous transition is observed, in agreement with the results found in the realistic quasi-2D model. For fixed values of $\Delta$ and $\alpha$, two timescales are at play, as seen in Fig. 3d. On the one hand, for a fixed value of $\tau_{s}$, the critical packing fraction increases with increasing drag coefficient $\gamma$. Indeed, higher drag implies that a denser system is needed to ensure sufficiently frequent collisions to sustain an active steady state. On the other hand, at a fixed value of $\gamma$, the critical packing fraction increases upon decreasing $\tau_{s}$, as this makes it easier to synchronize and hence lose energy. Another way to look at this is through the frequency of collision. When synchronization is at play in our system, we expect that at the critical packing fraction, the rate at which a particle synchronizes $1/\tau_{s}$ is of the same order of the frequency of collision which is itself, at a given temperature, an increasing function of $\phi$.
Figure 4: a, b, c) Power law behavior of observables for the realistic quasi-2D model with parameters $A=0.085\sigma$, $h=1.5077\sigma$ and $N=30000$ leading to $\phi_{c}\simeq 0.186$. d, e, f) Power law behavior of observables for the effective 2D model with parameters $\Delta/(\sigma\gamma)=0.15$, $\alpha=0.95$, $N=10^{6}$ leading to $\phi_{c}\simeq 0.1509$. The dashed lines represent the power law expected from a $2D$ system belonging to the conserved directed percolation universality class [tjhung2016criticality, odor2004universality]. a, d) Order parameter as a function of $\phi-\phi_{c}$. Expected exponent $\beta\simeq 0.64$ with $T^{ss}\sim(\phi-\phi_{c})^{\beta}$. b, e) Average over 100 runs of the evolution of the order parameter at $\phi_{c}$. Expected exponent $a\simeq 0.42$ with $T(t)\sim t^{-a}$. c, f) Time to relax to the steady state as a function of $|\phi-\phi_{c}|$. Expected exponent $\gamma\simeq 1.3$ with $\tau_{r}\sim|\phi-\phi_{c}|^{-\gamma}$.
@@ III Dynamical and structural properties of the phase transition
In this section, we examine the behavior of our system near the transition points. In particular, we highlight the differences between continuous and discontinuous transitions. Whenever possible, we also compare our findings with the behavior observed in similar systems reported in the literature.
@@ III.1 Nature of the transitions
Thus far, we have discussed the APT in terms of a discontinuous or continuous transition, simply based on whether there is an obvious jump in kinetic energy as a function of packing fraction. In this section, we take a closer look at the behavior around the transition point in order to examine the nature of the transition in more detail, for both the realistic quasi-2D and effective 2D model.
@@
For continuous transitions, in systems characterized by an infinite number of absorbing states and a conserved number of particles without additional symmetries, the associated phase transition is believed to belong to the conserved directed percolation universality class [PhysRevLett.89.190602, Ness2020, henkel2008non, lei2019hydrodynamics]. Evidence of this universality class, from the analysis of power law scaling, is given in Fig. 4 where simulation results for both the realistic quasi-2D and the effective 2D models are compared to the expected result for systems belonging to the conserved directed percolation universality class. In particular, panels a), b) and c) present data of the realistic quasi-2D model while d), e) and f) present those of the effective 2D model. In panels a) and d), we show that the order parameter variation, as a function of the distance from the critical point, obeys a power-law scaling characterized by an exponent very close to what is expected for conserved directed percolation universality class. In panels (b) and (e), we give the time evolution of the order parameter at the critical packing fraction $\phi_{c}$. For the effective 2D model, the expected scaling closely matches the evolution of the order parameter. For the realistic quasi-2D model, however, it is more challenging to discern the correct scaling due to the limited simulation time and the difficulty to accurately pinpoint the transition point. In panels d) and f), we provide data for the characteristic time $\tau_{r}$ to reach the steady state. Above $\phi_{c}$, $\tau_{r}$ is defined as the time at which the system first reaches its steady-state value within an arbitrary precision, which we set to 10 percent of the average steady-state temperature. These curves match well the expected scaling. Below $\phi_{c}$, $\tau_{r}$ is defined as the first time the system reaches an arbitrarily small value. We find significant variations in the curves depending on the chosen threshold, since the energy decreases as an exponential after the last collision. Here, we define the threshold as the lowest metastable steady-state temperature observed in a system that eventually decayed to the absorbing state. Note that the measured exponent is known to be dependent on the initial conditions [odor2004universality]. From the overall agreement with the expected scaling laws, we conclude that the phase transition is consistent with the conserved directed percolation universality class, although a definitive confirmation would require a finite size analysis.
@@
It is worth noting that our focus has been on systems exhibiting their phase transition at relatively high critical packing fraction ($\phi_{c}>0.1$). [lei2019hydrodynamics][lei2019hydrodynamics] Ref. lubeck2004universal. In contrast, when the transition occurs at very low $\phi_{c}$, the system may exhibit mean-field conserved directed percolation critical exponents due to effectively long-range interactions, as discussed in Refs. lei2019hydrodynamics, lubeck2004universal.
@@
Figure 5: Comparison of the order parameter evolution as a function of packing fraction in the realistic quasi-2D a) and effective 2D model b) for different system sizes. The main figures represent the evolution of the steady state temperature in the discontinuous case (we recall that the temperature is the xyxy-kinetic energy for the realistic quasi-2D model and the usual kinetic energy for the effective 2D model) for different system sizes and insets show the evolution of the order parameter of the last (the most dilute) active steady state ($\Delta T_{ss}$) as a function of the system size a) realistic quasi-2D model: $A=0.0626\sigma$ and $h=1.95\sigma$. b) Effective 2D model: $\Delta/\gamma\sigma=1.5$, $\alpha=0.95$, and $1/\gamma\tau_{s}=16.67$.
@@
Having clarified the nature of the continuous transition, we now turn to the discontinuous one and perform a finite size analysis to prove that the finite discontinuity does persist in the infinite system size limit. We define $\Delta T_{ss}$ which corresponds to the observed value of the order parameter of the most dilute system in the active state at a given system size and simulation time. For a continuous transition, we expect that at finite times, $\Delta T_{ss}$ decreases with $N$ and reaches 0 at infinite system size since the order parameter must vanish continuously. On the contrary, for discontinuous transition, $\Delta T_{ss}$ should not vanish. Evolution of the order parameter as a function of $\phi$ close to the discontinuous transition as well as $\Delta T_{ss}$ as a function of system size in inset are given, for the realistic quasi-2D and effective 2D models in the insets of Fig 5 a) and b) respectively. Starting from small system sizes, we see that increasing its size allows us to obtain a transition at a lower density, as would be expected as well for a continuous transition. However, beyond a given system size, $\Delta T_{ss}$ ceases to decrease and instead increases with size, indicating a discontinuous jump in the order parameter in the thermodynamic limit and confirming the first-order nature of the transition.
@@
We interpret such non-monotonic behavior of $\Delta T_{ss}$ as a nucleation dynamics – characteristic of discontinuous transitions – and better understood directly from the average time necessary for the system to reach an absorbing state at a given packing fraction in the metastable zone. We denote this time $t_{\mathrm{wait}}$ and study its behavior as a function of the number of particles for the effective 2D model in Fig. 6. The time $t_{\mathrm{wait}}$ acts as a proxy for the nucleation time, since the growth of the nucleus after it has reached its critical size should be fast compared to $t_{\mathrm{wait}}$. We only perform the analysis for the effective 2D model due to the prohibitive simulation time of the realistic quasi-2D model. For small system sizes, the waiting time increases with $N$ since the fluctuation required for the system to reach an absorbing state is of the order of the size of the system. Then, the waiting time reaches a maximum, which corresponds to a system size close to the size of the critical nucleus. It then decays due to the increase of available nucleation sites, until reaching a constant at larger system sizes due to the finite cooling rate imposed by $\gamma$ and the frequency of collision. Note that the decrease of the waiting time for system sizes larger than the critical nucleus is very sharp, while in an equilibrium we would expect a decay proportional to $1/N$. This might be the result of hydrodynamical instabilities or non-equilibrium structural properties at small $k$, inaccessible to small system sizes, which are common in granular systems [garzo2005instabilities, mitrano2011instabilities, mcnamara1994inelastic]. Note also that contrary to equilibrium systems, the nucleation dynamics, and especially the growth of the nucleus, exhibit an exotic behavior, both for the effective 2D and the realistic quasi-2D model. Since the order parameter is the activity or the temperature of the system, the nucleus cannot be stationary because, during its growth, active particles around it will quickly convert its inactive grains into active ones. We instead observe damped traveling waves of energy and density, slowly decreasing in amplitude until the absorbing state is reached (see Video 1 for the Realistic quasi-2D model and Video 2 for the effective 2D model). Moreover, contrary to Ref. neel2014dynamics we do not find a strongly diverging radius of the critical nucleus as we approach the stable fluid region but merely a strong increase of the effective energy barrier of nucleation. In the inset of Fig. 6 we confirm that the statistics of the waiting time are Poissonian when the reaching of the absorbing state is controlled by nucleation [khain2011hydrodynamics].
Interestingly, a very similar system has been studied by Lei, Hu, and Ni [lei2021barrier, lei2023does].lei2021barrier, lei2023does [di2016self, PhysRevResearch.2.013318].  neel2014dynamics They however found a different phenomenology. We provide additional discussions on this matter in Appendix. C.
@@
Figure 6: Waiting time in the effective 2D model for a system initialized in an active state to reach an absorbing state for different densities in the metastable zone; $\alpha=0.95$, $1/(\tau_{s}\gamma)=16.67$ and $\Delta/(\gamma\sigma)=1.5$. The inset shows the average waiting time as a function of its standard deviation. The dashed line is a linear scaling, showing that the process is Poissonian. Each point corresponds to an average over 50 independent runs.
@@ III.2 Structural properties of the transitions
Figure 7: Zoomed snapshots of the systems above ($\phi_{c}^{+}$, left) and below ($\phi_{c}^{-}$, right) the critical packing fraction for the realistic quasi-2D and effective 2D model. The particles with small velocities are drawn in black. The insets above $\phi_{c}$ represent the velocity distribution of particles corresponding to the snapshots in units of $\Delta$ for the effective 2D model and $\sigma f$ for the realistic quasi-2D model. The colors of the histograms of velocities match those of the particles in the snapshots (inactive are in black and active are in orange). The other insets represent the structure factor $\bm{S}_{\rho\rho}(k)$ of the corresponding snapshots as a function of wavenumber in units of $1/\sigma$.
Having clarified the nature of the transition, we now focus on dynamical and structural properties of the system close to it. In Fig. 7, we present, for both the effective 2D model and the realistic quasi-2D one, snapshots above and below $\phi_{c}$ for both type of transition, and the probability distribution of the velocities in the diffusive phase, and the long-wavelength behavior of the structure factor in the absorbing phase.
In the case of a continuous transition, the system reaches the absorbing state due to viscous drag forces. As commonly observed in similar situations, the critical point is characterized by avalanches of active particles that span the whole system [bak1991self, pruessner2013average, manna1990cascades]. This behavior is exemplified in the snapshots and leads to a distinctly identifiable population of inactive particles, discernible through the peak observed in the velocity distribution at the origin.
@@
Close to the transition point, $S_{\rho\rho}$ exhibits a scaling behavior of the form $S_{\rho\rho}(|\bm{k}|\rightarrow 0)\sim k^{\nu}$, with $\nu\simeq 0.5$ [wiese2024hyperuniformity]. The vanishing of the structure factor at $k=0$ is called hyperuniformity [torquato2018hyperuniform, torquato2016hyperuniformity, salvalaglio2024persistent] and is directly related to a suppression of density fluctuations on large length scales. This trend is consistent with the typical hyperuniform behavior observed at the critical point associated with the conserved directed percolation universality class [PhysRevLett.114.110602, lei2019hydrodynamics]. It has been observed that this scaling is expected to hold exactly at the transition point and to be lost below or above it [hexner2017enhanced]. In our models, however, system sizes are presumably not large enough for this hyperuniform scaling to break. Very recent works have successfully predicted hyperuniform scaling of the structure factor from a Renormalization Group approach by mapping the conserved directed percolation problem to the interface position at depinning for an elastic manifold [wiese2024hyperuniformity]. A direct analysis of the hyperuniformity from the Renormalization Group of the conserved directed percolation field theory was also performed in Ref. [ma2023theory] at first loop.
Finally, we report new hyperuniform scaling for the longitudinal and transverse velocity fields at criticality in Appendix. B.
@@
The velocity distributions in the active state exhibit peculiar scaling. As already noted, a finite portion of the system is almost completely inactive, hence there is a sharp peak at $v=0$, both for the realistic quasi-2D model and the effective 2D model to the velocity distribution. Moreover, at small velocity, we observe an exponential distribution with a nontrivial power: $\exp(-a|v|^{\delta})$ with $\delta\leq 1$, which then morphs into a simple exponential scaling: $\exp(-b|v|)$ at $v\simeq\Delta$. This effect is not surprising as non-Gaussian scalings are common in granular systems and, notably, it is known that the homogeneous cooling state exhibits a simple exponential tail while the homogeneously heated granular gas has an exponential tail with exponent $\delta=3/2$ [yu2020velocity, esipov1997granular, van1998velocity]. In particular, a simple exponential tail appears frequently [arnarson1998thermal, jarzynski1993universal] and is conjectured to be universal for systems with adiabatic energy change. However, this is not the case in our system, since the energy change during a collision is large compared to the energy of the particles. In our case, the pronounced non-Gaussian behavior is most likely due to the finite jump processes affecting the velocities, effectively transforming the dynamics into a generalized Ornstein-Uhlenbeck process with discrete jumps of size $\pm\Delta$. This makes the system analogous to a stochastic resetting problem, where an overdamped particle in a harmonic potential is reset away from the origin. Such processes are known to asymptotically yield stationary distributions with exponential or power-law tails [singh2020resetting, pal2015diffusion]. In Appendix A, we provide an explanation of this peculiar velocity distribution using an approximation for the Boltzmann equation.
@@
In the discontinuous case, the system primarily transitions into the absorbing phase in a manner similar to the cooling behavior observed in a granular gas. This is because most of the energy is dissipated through collisions between synchronized particles, with the drag playing a secondary role in this process. Interestingly, the presence of viscous friction is not a prerequisite for observing a transition. While it helps the system to reach sooner an absorbing state, the dissipative character of synchronized collisions alone is sufficient to deplete the entire energy of the system in the infinite time limit. This implies the formation of clusters similar to those observed in the inhomogeneous cooling state [goldhirsch1993clustering, garzo2005instabilities, brey1998hydrodynamics] for the realistic quasi-2D model, as can be seen in the snapshots and from the upturn of the structure factor at $k\rightarrow 0$. A scaling close to $k^{-3}$, but less steep is observed in the intermediate wavenumber region, which is roughly consistent with Porod’s law [bray1994theory, paul2017ballistic]. The inhomogeneous cooling is however arrested at a finite time due to the drag. In the effective 2D model, as collisions lead to a complete asynchronization of particles, the formation of clusters due to dissipative collisions is less prominent because high density regions of synchronized particles will quickly be shattered by any collision. However, even if the growth of the nucleus cannot proceed as the beginning of the inhomogeneous cooling in the effective 2D model, Porod’s law is approximately verified.
@@
The active state of the discontinuous transition does not exhibit a population of inactive particles but still has a marked non-Gaussian behavior, that, contrary to the case of the continuous transition, can be in part attributed to the dissipative collisions. We also provide the structure factor of the fluid close to the metastable region, which is in itself interesting. Lei and Ni [lei2023does] observed a hyperuniform scaling $S\sim k^{1.2}$ in a metastable fluid at very small wavenumbers. As already mentioned before, due to the dissipative collisions between synchronized particles, we instead observe a peak in the structure factor corresponding to the inverse average cluster size. We show in Appendix C that without dissipative collisions ($\alpha=1$) close to the metastable fluid, we recover a structure factor similar to the one found by Lei and Ni.